# Cascade Speculative Drafting for Even Faster LLM Inference

### 18 Dec 2023 v1

@@ Abstract
Speculative decoding enhances the efficiency of large language models (LLMs) by leveraging a draft model to draft for a larger target model to review. However, drafting in speculative decoding involves slow autoregressive generation and generating tokens of different importance with the same time allocation. These two inefficiencies lead to its suboptimal performance. To address this issue, we introduce Cascade Speculative Drafting (CS. Drafting), a novel approach that employs two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models. The Horizontal Cascade constitutes efficient time allocation in drafting with its optimality supported by our theoretical analysis. Combining both cascades, our CS. Drafting algorithm has achieved up to 72 percent additional speedup over speculative decoding in our experiments while keeping the same output distribution.
@@ 1 Introduction
The advent of Large Language Models (LLMs) like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but also found widespread application in user-interactive settings such as chatbots and virtual assistants. This application involves an extremely high number of users of up to hundreds of millions daily. To service in real-time at this scale, a low latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service making even a slight improvement over the latency of LLMs can contribute greatly to the community. Consequently, optimizing LLM latency has become a crucial and impactful research focus.
@@
Unfortunately, LLMs currently suffer from high latency especially for long-form generation as LLM can only generate one token at each run. The challenge is especially great during applications such as chatbots or virtual assistants due to the long generation of hundreds or even thousands of tokens. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step(with $k$ being a hyperparameter), and these tokens are then reviewed by a target model (which is larger and slower) in parallel. The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself. As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two factors: the expected number of tokens accepted by the target model per run, and the autoregressive efficiency of the draft model.
@@
Figure 1: In a single step of speculative decoding, the figure shows the probability of acceptance of draft tokens by their positions on GSM8K datasets, evaluated on FLAN-T5-small, base and large models. The draft model generates 30 tokens at each step.
@@
Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens. The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens that are proposed to and verified by LLMs at the next outer layer $\mathcal{M}_{d(i+1)}$. Once the verification process is completed, the Draft $\mathcal{M}_{d(i)}$ continues to generate new tokens given the accepted prefix tokens from previous layer $\mathcal{M}_{d(i-1)}$ as input.
Herein lies a trade-off: employing a larger draft model or generating more draft tokens per target run will improve the expected number of accepted tokens. However, either of these moves will increase the latency of the draft model significantly. As a result, the maximum speed improvement that the current speculative decoding algorithm can achieve is with an optimal balance between size and number of generated tokens. From previous work by Leviathan et al. (2023), it’s shown that very small models (less than 100 times the size of the target model) are optimal for most tasks experimented representing that the cost of drafting model can be the limiting factor of the speculative decoding algorithm. This observation leads to our central idea, due to the significant latency caused by the drafting model, improving drafting efficiency is the key to improving the performance of speculative decoding.
During our preliminary experiment, we found that speculative decoding with a simple non-neural language model can certainly improve the performance of a language model. This is because a simple non-neural language model is almost free of computational meaning that it’s free of the trade-off above. However, due to the limiting ability of such a simple non-neural language model, speculative decoding does not yield optimal results with a simple non-neural language model compared to a well-sized language model from the same model family of the target model. This makes a simple non-neural language model model suboptimal when a well-sized language model is available. We organize these observations into two laws of autoregressive generation of language models:
1. Without regard to generation length nor any additional hyperparameter tuning, a simple non-neural language model can always improve the generation speed of a language model by using speculative decoding over autoregressive generation.
2. When there is a well-sized smaller model(of the same architecture, distilled from the target model or trained from a similar corpus) and a task with enough expected generation length, the smaller model can be used to improve the generation speed of the target model over the simple non-neural language model.
@@
These two laws provide us with insight into designing a new algorithm for even further speed improvement. We propose a new algorithm Cascade Speculative Drafting (CS. Drafting) a speculative execution algorithm with a paradigm that radically excludes any autoregressive generation from neural language models and exhaustively substitutes any simple non-neural language models with superior neural language models. The resulting algorithm contains multiple draft models with the smallest being a simple non-neural language model, with each neural draft model reviewing generations from a smaller model while proposing its reviewed content to a larger draft model or the target model. We name this layered speculative decoding design the Vertical Cascade.
@@
Moreover, our experimental findings reveal an intriguing pattern: in a drafting step, the tokens generated later in the sequence by the draft model exhibit a progressively lower probability of being accepted by the target model, as illustrated in Figure 1. (Note that this is different from the acceptance rate which is the probability of the token being accepted conditioned on the previous token being accepted) The observation suggests that the later tokens generated by draft models, with a higher likelihood of rejection, contribute much less to the expected number of accepted tokens while creating the same latency. To address this, we propose using smaller, faster draft models $\mathcal{M}_{d(i-1)}$ for generating these high-rejection tokens, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS. Drafting approach illustrated in Figure 2. Integrating these vertical and horizontal cascades, our method innovatively mitigates LLM latency issues.
We evaluated our proposed CS. Drafting on a variety of tasks and model sizes. The experimental results demonstrate the superiority of our method. Moreover, theoretical analysis is provided to prove the optimality of our proposed method. Ablation studies further verify the effectiveness of our method.
In summary, our main contributions are as follows:
@@
  • We introduce a novel algorithm, Cascade Speculative Drafting (CS. Drafting), a speculative-execution-based algorithm improving language models inference speed without compromising on generality or overall efficiency.
  • Our work includes a theoretical analysis affirming the optimality of our CS. Drafting approach.
  • We have established a new evaluation method, standardized walltime improvement, tailored for algorithms based on speculative decoding to ensure reproducibility and facilitate consistent comparisons across different studies.
  • Through our empirical study, we demonstrate that our CS. Drafting algorithm consistently achieves lower latency compared to vanilla speculative decoding across different tasks and settings. With a speedup of up to 72% over speculative decoding, this empirical evidence underscores the practical benefits and efficiency gains of our proposed approach.
@@ 2.1 Speculative Decoding Algorithm
The core concept of speculative decoding involves a novel algorithm that accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models. generation
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{1},...,x_{k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively. In the reviewing stage, let $x_{i}\in\{x_{1},...,x_{k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$ ; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ an then reject any token following $x_{i}$. Such a design guarantees the output is the same as sampling autoregressively using the target model alone.
This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs. Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of autoregressive models such as Transformers (Leviathan et al., 2023).
2.2 Notation and Definition
We introduce notions and definitions that are used throughout the paper. Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token. In the simpler case of the speculative decoding, $\mathcal{M}_{i}=q$ for any $i$. Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
Cost coefficient $c_{i}$ is the ratio of time for a single run of $\mathcal{M}_{i}$ over $\mathcal{M}_{t}$.
Lenience in speculative decoding is a parameter to loosen the review process by the target model for further speedup with the tradeoff that the generation result might not be the same as the target model. Yet in a recursive speculative decoding setup, lenience in the inner speculative decoding will not affect the final output. With lenience $l\in[1,\infty)$, The acceptance condition of $x_{i}$ speculative decoding will become $\mathcal{M}_{d}(x)[i]<l*\mathcal{M}_{t}(x)[i]$; otherwise, with a probability of $1-\frac{l*\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$ , we reject the $x_{i}$ and any following tokens.
Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on speculative decoding over the vanilla generation of the target model.
@@ 3 Cascaded Speculative Drafting
  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ []
   $\triangleright$ Unpack the current parameters from
@@
   $\triangleright$ Base case when the MaG generate
@@
     curK $\leftarrow$ the submatrix of $K_{nn}$ from with left-top cornor (i + 1, j) to the right-most bottom
     while length of curGen is less than $k_{i}$ do
        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, curK, $l$)
        curGen $\leftarrow[x_{1},..,x_{s}]$
@@
  return $prefix+[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
@@ 3.1 Horizontal Cascade
One key observation that leads to our experiment is that during the drafting steps of speculative decoding, not all drafting tokens are made equal, as the first draft token is exponentially easier to be accepted than the $k$-th draft token since the expected acceptance rate of the $q_{i}$ decreases exponentially as $i$ increases. In Corollary 4.3, we proved the exponential decrease above and also observed that the derivative of EWIF with respect to $\alpha_{1}$ is also exponentially greater than the derivative of EWIF with respect to $\alpha_{k}$ representing the effort in improving model acceptance rate on the first token, though having the same cost, is exponentially more valuable than the effort on the last token. As a result, given multiple models available for drafting, we propose cascade speculative decoding which involves using a larger model to generate the first few tokens and a smaller model to generate the remaining tokens.
In addition, assuming $\mathcal{M}_{d1},\mathcal{M}_{d2}$ being two draft models [^1] The notation of $\mathcal{M}_{d1}$ is different from $\mathcal{M}_{i}$ When the $d$ appeared in the subscript, it means the model can be used for generating multiple tokens. with $\mathcal{M}_{d1}$ having high cost and acceptance rate then $\mathcal{M}_{d2}$ during drafting of $i$-th token with cascade decoding, a more general setup can be achieved by using an interpolated draft model by assigning a probability $f(i)$ such that with probability $f(i)$ the $i$-th token is generated by $\mathcal{M}_{d1}$; otherwise, it’s generated by $\mathcal{M}_{d2}$. This interpolated draft model setup can provide a smoother transition of cost and acceptance rate. However, we prove that such a setup will not improve the walltime improvement factor in theorem 4.4.
3.2 Vertical Cascade
With our proven results of theorem 4.4 that for optimal cascade sampling of two draft model $\mathcal{M}_{d1},\mathcal{M}_{d2}$, each draft token will be generated by one of the draft models, we observe one additional inefficiency even with the already improved efficiency with horizontal cascade. As we have two draft models $\mathcal{M}_{d1},\mathcal{M}_{d2}$ with $\mathcal{M}_{d1}$ having better acceptance rate and higher time cost, we hypothesize that another round of speculative sampling replacing autoregressive generation of $\mathcal{M}_{d1}$ for drafting can further improve the efficiency of cascade sampling. Furthermore, leniency introduced by Leviathan et al. (Leviathan et al., 2023) provides a tradeoff between quality and speed. In practice, such a quality drop can be undesirable; however, the Cascade Speculative Drafting algorithm does not have such a limitation. Since the final quality is always guaranteed by the target model, any lenience in speculative decoding between the draft models will not affect the final output. In addition, we found that lenience can potentially bring further speedup as the optimal matching to the larger draft model in speculative drafting might not reduce the generation speed as much as the speedup introduced by lenience.
3.3 Max-gram for Better simple non-neural language model
A simple non-neural language model is special to speculative decoding as it almost incurs no cost when drafting for its target model. This means despite a slight implementation overhead, it can almost always replace autoregressive generation using greedy decoding, sampling, or beam search. In addition, in the CS Drafting algorithm, a simple non-neural language model also generates for the draft model that reviews the most tokens making the output alignment between the simple non-neural language model and the smallest draft model very important for the speedup of our algorithm.
During our pursuing of a better simple non-neural language model, we noticed a general pattern. When language models generate, some words and phrases in the input query appear repeatedly in the generation result. A common sense that the answer should have words and phrases overlapping with the question.
Based on the observation, we designed the Max-Gram(MaG) algorithm that greedily finds maximal matches in the initial input(or existing generation) and tokens from the end of the generation, and in case there is no match, we use a bigram model pretrained on the Wikipedia. (We chose Wikipedia to not affect the generality of our experiment) To our surprise, despite its simplicity, we found this simple and general algorithm very effective in aligning with language model model output. We included a GPU-friendly version of the Max-Gram algorithm in the appendix A.

### 21 Dec 2023 v2

@@ Abstract
Speculative decoding reduces the latency of large language model (LLM) inference by leveraging a draft model to draft for a larger target model to review. However, drafting in speculative decoding involves slow autoregressive generation and allocates the same amount of time to generating tokens of varying importance. These two inefficiencies lead to its suboptimal performance. To address this issue, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade constitutes efficient time allocation in drafting. Combining both cascades, our CS Drafting algorithm has achieved up to a 72 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model.
@@ 1 Introduction
The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but have also found widespread application in user-interactive settings such as chatbots and virtual assistants. This application involves an extremely high number of users, up to hundreds of millions daily. To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the provider and the community. Consequently, optimizing LLM latency has become a crucial and impactful research focus.
@@
Unfortunately, LLMs currently suffer from high latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter), and these tokens are then reviewed by a target model (which is larger and slower) in parallel. The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself. As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two main factors: the expected number of tokens accepted by the target model per run, and the latency of drafting.
@@
Figure 1: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models. The draft model generates 30 tokens at each step.
@@
Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens. The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens proposed to and verified by LLMs at the next outer layer, $\mathcal{M}_{d(i+1)}$. This recursive process forms the vertical cascade. After a vertical cascade, a smaller $\mathcal{M}_{d(i)}$ will serve as the target model and perform a vertical cascade. The process repeats until the smaller model becomes a statistical language model that will generate autoregressively forming the horizontal cascade.
[…]
In light of this, there exist a type of draft models, including statistical language models like bigram language models, that are immune to drafting efficiency issues. Owing to their simple structure, these models incur negligible latency and computational resource costs compared to neural language models, thereby avoiding the aforementioned trade-off. Therefore, assuming enough parallel computing power to perform computation on additional $k$ tokens with similar time cost (Leviathan et al., 2023), utilizing speculative decoding with a statistical language model as the draft model can generally and intuitively enhance the efficiency of a neural language model. However, due to their limited capabilities, speculative decoding with statistical language models alone may not yield optimal results compared to using a well-sized neural language model from the same family as the target model when the neural language model is available.
@@
Based on the above analysis, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model. Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model. We refer to this tiered speculative decoding approach as the Vertical Cascade.
@@
Furthermore, our analysis in Figure 1 reveals an interesting pattern in the expected acceptance rate during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model [^1] A theoretical analysis is provided in Section 4. Note that this probability differs from the acceptance rate, which is the probability of a token being accepted conditioned on the previous tokens’ acceptance.. This indicates that later tokens from draft models, more prone to rejection, contribute less to the expected number of accepted tokens, yet incur the same latency. Based on this observation and our practice, we suggest the use of smaller, faster draft models for generating these high-rejection tokens, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 2.
Through theoretical analysis and empirical studies, we demonstrate that the CS Drafting algorithm outperforms speculative decoding in terms of latency across various tasks and settings, achieving an additional speedup of up to 72% over speculative decoding. These findings highlight the practical advantages and efficiency enhancements offered by both vertical and horizontal cascades.
Our main contributions are summarized as follows:
@@
  • We introduce Cascade Speculative Drafting (CS Drafting), a speculative-execution-based algorithm that improves language model inference speed without sacrificing generation quality.
  • We provide theoretical analyses supporting the effectiveness of the proposed CS Drafting approach.
  • We conduct empirical experiments showing that CS Drafting achieves further speedup over speculative decoding across different tasks and settings.
@@ 2.1 Speculative Decoding Algorithm
The core concept of speculative decoding involves a novel algorithm that accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$ In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$. At the end, the target model will generate one additional token following the accepted tokens. Such a design guarantees the output is the same as sampling autoregressively using the target model alone (Leviathan et al., 2023).
This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs. Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of LLMs.
@@ 3 Cascaded Speculative Drafting
   $\triangleright$ Initialize curGen and curProb
  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ a list of ones of length prefix
   $\triangleright$ Unpack the a list of $k$ for current function call
@@
   $\triangleright$ Base case of the recursive call when the MaG generate
@@
     curK $\leftarrow$ the submatrix of $K_{nn}$ from with the top-left corner at $(i+1,j)$ extending to the bottom-right corner
     while length of curGen - prefix.length is less than $k_{i}$ do
        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, False, curK, $l$)
        curGen $\leftarrow[x_{1},..,x_{r}]$
@@
  return $[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
3.1 Vertical Cascade
The speculative decoding algorithm relies on the autoregressive generation of a smaller draft model. However, despite its smaller size, the draft model must run $k$ times for each target model run, and the cost can still be significant. To further reduce latency in the drafting process, we propose using an even smaller model to assist in drafting and employing the original draft model to review the generation of this smaller model. This process can be done recursively until it reaches the smallest draft model. The smallest draft model can be a statistical language model, such as a bigram language model, where the cost is negligible compared to neural language models. We refer to this recursive process as Vertical Cascade.
Furthermore, to further improve speed, we can introduce leniency, a hyperparameter in speculative decoding that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023). With leniency $l\in[1,\infty)$, the acceptance condition for token $x_{i}$ in speculative decoding becomes $\mathcal{M}{d}(x)[i]<l\times\mathcal{M}{t}(x)[i]$. Otherwise, with a probability of $1-\frac{l\times\mathcal{M}{t}(x)}{\mathcal{M}{d}(x)}$, we reject $x_{i}$ and any following tokens. While the reduced quality introduced by leniency in speculative decoding is generally undesirable in practice, it fortunately does not affect the final output in our vertical cascade approach—we can only apply leniency in the inner speculative decoding. As long as leniency is not applied to the target model, the final quality is always guaranteed by the target model.
@@ 3.2 Horizontal Cascade
As illustrated in Figure 1, another key observation is that during the drafting steps of speculative decoding, not all drafting tokens are created equal. The first draft token is more likely to be accepted as it only depends on itself; the last token is rarely accepted, as it has a chance of being reviewed only if all preceding tokens are accepted. From another perspective, one could also say that the first draft token bears the responsibility for all succeeding tokens, while the last token bears none. In Section 4, we provide a theoretical analysis of this phenomenon.
Based on this observation, we propose better allocation of the time cost of drafting by using larger models to generate the preceding draft tokens and smaller models for the succeeding tokens. This design allows more time to be allocated to more effective tokens, leading to a reduction in overall latency. We refer to this design as the Horizontal Cascade.
3.3 Max-gram for Better Statistical Drafting
A statistical language model is special to speculative decoding, as it incurs almost no cost when drafting for its target model. In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content. Based on this observation, we designed the Max-Gram (MaG) algorithm. It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation. In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality of our experiment). Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs. We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.